OpenStudy (anonymous):

Simplify: (((6-3x)^1/2)+((3/2)x((6-3x)^(-1/2))))/(6-3x)

OpenStudy (zzr0ck3r):

\[\frac{\sqrt{6-3x}+\frac{3x}{2\sqrt{6-3x}}}{6-3x}\]?

OpenStudy (theeric):

@r2sahu , have you given it a shot?

OpenStudy (zzr0ck3r):

@r2sahu is this what you need simplified?

OpenStudy (anonymous):